# Myopically Verifiable Probabilistic Certificates for Safe Control and Learning

### v1

@@ V-B2 Long-term avoidance and finite-time reachability
$\mathbf{F}_{k}(X):=$ $\sup_{N\in\mathcal{N}_{d}}\mathbb{P}\{X_{\tau}\in{\mathcal{C}}(L),\forall\tau\in\{k,\cdots,T_{d}\}\mid X_{k}=X\}$
@@
::=:= $\sup_{N\in\mathcal{N}_{d}}\mathbb{P}\{X_{\tau}\in{\mathcal{C}}(L)\text{ for some }\tau\in\{k,\cdots,T_{d}\}\mid X_{k}=X\}$
@@
in case of finite-time reachability. Here, ${\mathcal{C}}^{c}$ is the complement of the safe set, i.e., unsafe set, and the expectations are taken assuming $U_{\tau}=N(X_{\tau}),\forall\tau\in\{k,\cdots,T_{d}\}$. Note that the right hand side of both (74) and (75) are special cases of the problem defined in [12]. The goal is to estimate $\mathbf{F}_{k}(X)$ for $X\in\mathcal{X}$ and $k=\{T_{d},\cdots,0\}$ backward in time using DP.
@@
is the Gaussian function. The constants $c_{i,j}$, $i\in\{1,\cdots,M\}$, $j\in\{1,\cdots,n\}$ are sampled from a uniform distribution in ${\mathcal{C}}$ (or a bounded subset of interest in ${\mathcal{C}}$) for long-term avoidance and in ${\mathcal{C}}^{c}$ (or a bounded subset of interest in ${\mathcal{C}}^{c}$) for finite-time reachability. The constants $\nu_{i,j}$, $i\in\{1,\cdots,M\}$, $j\in\{1,\cdots,n\}$ are randomly sampled on $\mathbb{R}^{+}$. We solve the weight $w_{i}$ for the linear program (78) to approximate $\hat{\mathbf{F}}_{k}(X)$. The entire recursive approximation process is summarized in Algorithm 2. The output $\hat{\mathbf{F}}_{0}(X)$ is an approximation of $\mathbf{F}(Z)$ in terms of long-term avoidance and finite-time reachability.
@@
V-C Safe Control Algorithms
Here, we propose two safe control schemes based on the safety conditions introduced in subsection V-A. In both schemes, the value of $\mathbf{F}$ is defined in (25) as type 1 or 2 when the safety specification is given as forward invariance condition, and as type 3 or 4 when the safety specification is given as forward convergence condition.
V-C1 Additive modification
@@
$N^{\prime}(X_{t},L_{t},T_{t})=N(X_{t})+\kappa(Z_{t})({\mathcal{L}}_{\tilde{g}}\mathbf{F}(Z_{t}))^{\intercal}.$ (80)
@@
$A\mathbf{F}(Z_{t})$ (81)
$=$ ${\mathcal{L}}_{\tilde{f}}\mathbf{F}(Z_{t})+({\mathcal{L}}_{\tilde{g}}\mathbf{F}(Z_{t}))N(X_{t})+\kappa{\mathcal{L}}_{\tilde{g}}\mathbf{F}(Z_{t})\left({\mathcal{L}}_{\tilde{g}}\mathbf{F}(Z_{t})\right)^{\intercal}$
$\qquad+\frac{1}{2}\text{tr}\left(\tilde{\sigma}\tilde{\sigma}^{\intercal}\operatorname{Hess}{\mathbf{F}}(Z_{t})\right).$ (82)
As $\kappa$ is non-negative, the term $\kappa{\mathcal{L}}_{\tilde{g}}\mathbf{F}\left({\mathcal{L}}_{\tilde{g}}\mathbf{F}\right)^{\intercal}$ in (81) takes non-negative values. This implies that the second term additively modify the nominal controller output $N(X_{t})$ in the ascending direction of the safety probability.
@@
$\kappa^{*}=&&\operatorname*{arg\,min}_{\kappa}&\ \ \mathbb{E}\left[J(\kappa,Z_{t})\right]\\ &&\text{s.t. }\ \ \mathbb{E}&[D_{\mathbf{F}}(Z_{t},N^{\prime}(X_{t},L_{t},T_{t}))\\ &&&+\alpha\left(\mathbf{F}(Z_{t})-(1-\epsilon)\right)]\geq 0,\end{aligned}$ (83)
where $J(\kappa,Z_{t})$ is the objective function of choice to minimize. The expectations in (83) are taken over the process of the closed-loop system with (80).
V-C2 Constrained optimization
@@
[…]
@@
where $J$ is the objective function of MPC, $\bar{H}$ is the outlook prediction horizon of MPC. The control action $N^{\prime}(X_{t},L_{t},T_{t})$ takes the value of $U_{t}$ in $\mathbf{u}$. The optimization problem is designed to satisfy the assumptions of Theorem 1 to comply with the safety specification (26). Note that once we formulate the safe control problem into constrained optimization, one can add additional constraints into the same optimization problem [53, 54]. It is worth noting that the MPC prediction horizon $\bar{H}$ can be shorter than the horizon to ensure safety because the safety constraint is only imposed for $Z_{t}$ in (84). In the special case where the MPC prediction horizon is 1, the constrained optimization becomes
@@
where $J:\mathbb{R}^{m}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$ is the objective function to be minimized. The constraint of (85) imposes that (39) holds at all time $t$, and can additionally capture other design restrictions. [^11] For example, $N^{\prime}$ is Lipschitz continuous when $J(N(x),u)=u^{\intercal}H(x)u$ with $H(x)$ being a positive definite matrix (pointwise in $x$).
Both additive modification and constrained optimization are commonly used in the safe control of deterministic systems (see [44, subsection II-B] and references therein). These existing methods are designed to find control actions so that the vector field of the state does not point outside of the safe set around its boundary. In other words, the value of the barrier function will be non-decreasing in the infinitesimal future outlook time horizon whenever the state is close to the boundary of the safe set. However, such myopic decision-making may not account for the fact that different directions of the tangent cone of the safe set may lead to vastly different long-term safety. In contrast, the proposed control policies (80) and (85) account for the long-term safe probability in $\mathbf{F}$, and are guaranteed to steer the state toward the direction with non-decreasing long-term safe probability when the tolerable long-term unsafe probability is about to be violated. Note that $\mathbf{F}$ can be computed offline and the controller only needs to myopically evaluate (80) or solve (85) in real-time execution. In this case, the online computation efficiency is comparable to common myopic barrier function-based methods in a deterministic system.
The overall safe control strategy is shown in Algorithm 3. Note that we present the algorithm in discrete time as most digital controller requires discretization. We first initialize $H$ in (28) for outlook time horizon, time step $\Delta t$, initial state $X_{0}$, nominal controller $N$, and maximum simulation time $T_{\text{max}}$. Then we acquire the safety probability of each state using the computation methods introduced in section V-B. At each time step $t$, we use the safe control methods in section V-C to calculate the safe control action $U_{t}$. We execute this control action to step the dynamics and this together forms our overall safe control strategy.
@@
Initialize $H$, $\Delta t$, $X_{0}$, $N$, $T_{\text{max}}$
Get $\mathbf{F}(X)$ for $X\in\mathcal{X}$ through Algorithm 1 or Algorithm 2
@@
VI Integration with Reinforcement Learning
In this section, we discuss how to integrate our proposed safe control method with reinforcement learning methods for control problems. We follow the standard setups for reinforcement learning as described in [55] with discrete time, and use $\{X_{t}\}$ and $\{U_{t}\}$ for state and action instead for notation consistency.
@@ VI Integration with Reinforcement Learning
where $\text{Proj}_{(\cdot)}$ is operator of taking projection onto ()($\cdot$)( ⋅ ), and $\mathcal{S}:=\{U:D_{\mathbf{F}}(Z_{t},U)\geq-\alpha\left(\mathbf{F}(Z_{t})-(1-\epsilon)\right)\}$ is the set of safe control actions. We represent this process by
@@
VI-A Integration with Policy Gradient
@@
which is the expected cumulative rewards under policy $\pi_{\theta}$.
@@
[…]
@@
[…]
where we know that {dmath} ∇logθτ) = ∇log[(X0) ∏t=0Hr-1πθ(^Ut∣Xt) G(Ut∣Xt, ^Ut) (Xt+1∣Xt, Ut)] = ∇[log((X0)) + ∑t=0Hr-1(log(πθ(^Ut∣Xt)) + log(G(Ut∣Xt, ^Ut)) + log((Xt+1∣Xt, Ut))) ] = ∇[∑t=0Hr-1log(πθ(^Ut∣Xt)) ]= ∑t=0Hr-1∇log(πθ(^Ut∣Xt)). The third equation holds because $G$ is independent from the policy $\pi_{\theta}$. Combining (95) and (95), we get (92). ∎
With (92), one can perform policy optimization $\theta\leftarrow\theta+\eta\nabla V^{\theta}$ to train the reinforcement learning agent with $\eta$ being the learning rate [55]. We conclude the procedures to apply the proposed safety filter to policy gradient in Algorithm 4. For each policy gradient iteration in the total $N_{\text{iter}}$ iterations, we collect sample trajectories for $N_{\text{eps}}$ episodes, and calculate the modified policy gradient using (92). Then we can perform policy optimization with the modified policy gradient.
@@
for $i=1:N_{\text{iter}}$ do
     for $k=1:N_{\text{eps}}$ do
@@
     Estimate $\nabla V^{\theta_{i}}$ in (92) with $R(\tau_{k}),k\in\{1,\cdots,N_{eps}\}$
@@
VI-B Integration with Q-learning
@@
::=:= $r(x,u)+\gamma\;{\mathbb{E}}_{{x^{\prime}\sim\mathbb{P}(X_{t+1}\mid X_{t}=x,U_{t}=u)}}\left[\max_{u^{\prime}\in\mathcal{U}}Q\left(x^{\prime},u^{\prime}\right)\right].$
Then, we have the Q-learning update rule for each iteration $i$ being
@@
where $\eta_{i}$ is the learning rate. From [56], when the reward function $r$ is bounded, the learning rate satisfies $10\leq\eta_{i}<$ 1, and
@@
the Q-learning converges to the optimal Q function asymptocially with probability 1.
The procedures for the proposed safe Q-learning method is summarized in Algorithm 5. Note that the algorithm only learns Q function. There are many choices for how to obtain the control policy using Q function (e.g., $\epsilon$-greedy).

### v2

@@ 5.2.2 Long-term avoidance and finite-time reachability
$\mathbf{F}_{k}(X):=$ $\sup_{N\in\mathcal{N}_{d}}\mathbb{P}\{X_{\tau}\in{\mathcal{C}}(L),$
$\quad\forall\tau\in\{k,k+1,k+2\cdots,T_{d}\}\mid X_{k}=X\}$
@@
$:=$ $\sup_{N\in\mathcal{N}_{d}}\mathbb{P}\{X_{\tau}\in{\mathcal{C}}(L)$
 $\quad\text{ for some }\tau\in\{k,k+1,k+2\cdots,T_{d}\}\mid X_{k}=X\}$
@@
in case of finite-time reachability. Here, ${\mathcal{C}}^{c}$ is the complement of the safe set, i.e., unsafe set, and the expectations are taken assuming $U_{\tau}=N(X_{\tau}),\forall\tau\in\{k,k+1,k+2\cdots,T_{d}\}$. Note that the right hand side of both (58) and (59) are special cases of the problem defined in [13]. The goal is to estimate $\mathbf{F}_{k}(X)$ for $X\in\mathcal{X}$ and $k=\{T_{d},T_{d}-1,T_{d}-2\cdots,0\}$ using DP. Note that the indexing is in reverse time order as we solve for $\mathbf{F}_{k}(X)$ backward in time.
@@
is the Gaussian function. The constants $c_{i,j}$, $i\in\{1,2,3,\cdots,M\}$, $j\in\{1,2,3,\cdots,n\}$ are sampled from a uniform distribution in $\mathcal{C}$} (or a bounded subset of interest in $\mathcal{C}$}) for long-term avoidance and in ${\mathcal{C}}^{c}$ (or a bounded subset of interest in ${\mathcal{C}}^{c}$) for finite-time reachability. The constants $\nu_{i,j}$, $i\in\{1,2,3\cdots,M\}$, $j\in\{1,2,3,\cdots,n\}$ are randomly sampled on $\mathbb{R}^{+}$. [^10] In practice, we can uniformly sample $v_{i,j}\sim\text{Unif}(1,2)$. As the Gaussian Radial basis functions are universal approximators with sufficiently many basis elements [63], $\hat{\mathbf{F}}_{0}(X)$ can approximate the true value with arbitrary precision as the number of basis elements ($M$ and $n$) increase. We solve the weight $w_{i}$ for the linear program (62) to approximate $\hat{\mathbf{F}}_{k}(X)$. The entire recursive approximation process is summarized in Algorithm 2. The output $\hat{\mathbf{F}}_{0}(X)$ is an approximation of $\mathbf{F}(Z)$ in terms of long-term avoidance and finite-time reachability.
@@
Remark 3.
Based on [58, 64], with a large enough number of samples, the estimation error can be made arbitrarily small with a sufficient number of samples. Besides, the long-term safety probability of Type 1 and Type 3 are solutions to partial differential equations (PDEs) [65, 66]. There exist numerical methods to solve these PDEs such as the Crank–Nicolson method [67] that are known to be convergent, i.e., the numerical solution approaches the exact solution as the grid size approaches 0.
5.3 Safe Control Algorithms
Here, we propose two safe control schemes based on the safety conditions introduced in subsection 5.1. In both schemes, the value of $\mathbf{F}$ is defined in (13) as type 1 or 2 when the safety specification is given as forward invariance condition, and as type 3 or 4 when the safety specification is given as forward convergence condition.
5.3.1 Additive modification
@@
$N^{\prime}(X_{t},L_{t},T_{t})=N(X_{t})+\kappa(Z_{t})({\mathcal{L}}_{\tilde{g}}\mathbf{F}(Z_{t}))^{\top}.$ (64)
@@
$D_{\mathbf{F}}(Z_{t},N^{\prime}(X_{t},L_{t},T_{t}))={\mathcal{L}}_{\tilde{f}}\mathbf{F}(Z_{t})+({\mathcal{L}}_{\tilde{g}}\mathbf{F}(Z_{t}))N(X_{t})$
$\quad+\kappa{\mathcal{L}}_{\tilde{g}}\mathbf{F}(Z_{t})\left({\mathcal{L}}_{\tilde{g}}\mathbf{F}(Z_{t})\right)^{\top}+\frac{1}{2}\text{tr}\left(\tilde{\sigma}\tilde{\sigma}^{\top}\operatorname{Hess}{\mathbf{F}}(Z_{t})\right).$ (65)
As $\kappa$ is non-negative, the term $\kappa{\mathcal{L}}_{\tilde{g}}\mathbf{F}\left({\mathcal{L}}_{\tilde{g}}\mathbf{F}\right)^{\top}$ in (5.3.1) takes non-negative values. This implies that the second term additively modify the nominal controller output $N(X_{t})$ in the ascending direction of the safety probability.
@@
$\begin{aligned} \kappa^{*}=&&\operatorname*{arg\,min}_{\kappa}&\ \ \mathbb{E}\left[J(\kappa,Z_{t})\right]\\ &&\text{s.t. }\ \ &D_{\mathbf{F}}(Z_{t},N^{\prime}(X_{t},L_{t},T_{t}))\\ &&&+\alpha\left(\mathbf{F}(Z_{t})-(1-\epsilon)\right)\geq 0,\end{aligned}$ (66)
where $J(\kappa,Z_{t})$ is the objective function of choice to minimize. The expectations in (66) are taken over the process of the closed-loop system with (64). Note that (66) only needs to be solved once offline, then the safe control can be calculated via (64) online. In practice, through polynomial parameterization (or other basis functions parameterization) of $\kappa$, (66) can be effectively solved with sum-of-squares optimizations [69, 70], similar to what has been done in the barrier certificate literature [26].
5.3.2 Constrained optimization
@@
$\qquad\text{s.t.}\ \ D_{\mathbf{F}}(Z_{t},U_{t})\geq-\alpha\left(\mathbf{F}(Z_{t})-(1-\epsilon)\right),$
@@
where $J$ is the objective function of MPC, $\bar{H}$ is the outlook prediction horizon of MPC, and $D_{\mathbf{F}}(Z_{t},U_{t})$ is the infinitesimal generator defined in (26). The control action $N^{\prime}(X_{t},L_{t},T_{t})$ takes the value of $U_{t}$ in $\mathbf{u}$. The optimization problem is designed to satisfy the assumptions of Theorem 1 to comply with the safety specification (14). Note that once we formulate the safe control problem into constrained optimization, one can add additional constraints into the same optimization problem [71, 72]. It is worth noting that the MPC prediction horizon $\bar{H}$ can be shorter than the horizon to ensure safety because the safety constraint is only imposed for $Z_{t}$ in (67). In the special case where the MPC prediction horizon is 1, the constrained optimization becomes
@@
where $J:\mathbb{R}^{m}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$ is the objective function to be minimized. Note that one can use a performance-oriented nominal controller to define the objective function where $J$ penalizes the deviation from the nominal controller. The constraint of (68) imposes that (27) holds at all time $t$, and can additionally capture other design restrictions. [^11] For example, $N^{\prime}$ is Lipschitz continuous when $J(N(x),u)=u^{\top}H(x)u$ with $H(x)$ being a positive definite matrix (pointwise in $x$).
Both additive modification and constrained optimization are commonly used in the safe control of deterministic systems (see [65, subsection II-B] and references therein). These existing methods are designed to find control actions so that the vector field of the state does not point outside of the safe set around its boundary. In other words, the value of the barrier function will be non-decreasing in the infinitesimal future outlook time horizon whenever the state is close to the boundary of the safe set. However, such myopic decision-making may not account for the fact that different directions of the tangent cone of the safe set may lead to vastly different long-term safety. In contrast, the proposed control policies (64) and (68) account for the long-term safe probability in $\mathbf{F}$, and are guaranteed to steer the state toward the direction with non-decreasing long-term safe probability when the tolerable long-term unsafe probability is about to be violated. Note that $\mathbf{F}$ can be computed offline and the controller only needs to myopically evaluate (64) or solve (68) in real-time execution. In this case, the online computation efficiency is comparable to common myopic barrier function-based methods in a deterministic system.
The overall safe control strategy is shown in Algorithm 3. Note that we present the algorithm in discrete time as most digital controller requires discretization. We first initialize $H$ in (16) for outlook time horizon, time step $\Delta t$, initial state $X_{0}$, nominal controller $N$, and maximum simulation time $T_{\text{max}}$. Then we acquire the safety probability of each state using the computation methods introduced in section 5.2. At each time step $t$, we use the safe control methods in section 5.3 to calculate the safe control action $U_{t}$. We execute this control action to step the dynamics and this together forms our overall safe control strategy.
@@
Initialize $H$, $\Delta t$, $X_{0}$, $N$, $T_{\text{max}}$ Get $\mathbf{F}(X)$ for $X\in\mathcal{X}$ through Algorithm 1 or Algorithm 2
@@
So far, we consider safe control problems and aim to find a constraint on action to be imposed on some nominal controller or in model predictive control. In the following section, we consider reinforcement learning and we aim to learn an optimal controller for a given objective function that also satisfies the safety constraints during and after learning.
6 Integration with Reinforcement Learning
In this section, we discuss how to integrate our proposed safe control method with reinforcement learning (RL) methods for control problems. Specifically, we show how to optimize the control policy with RL in a way that accounts for the potential modification from the safety certificate, to meet RL objectives while ensuring long-term safety. In the proposed RL methods that incorporate the proposed safety constraints, its performance guarantee (Theorem 1) also applies.
Specifically, we consider policy gradient and Q-learning, as they are the two major approaches for solving reinforcement learning problems [73, 74]. We follow the discrete-time setup for reinforcement learning as described in [75, Chapter 3], and use $\{X_{t}\}$ and $\{U_{t}\}$ for state and action instead for notation consistency. Note that in this setup, the ability to perform roll-outs of the system dynamics is assumed, which is essential for calculating values such as safety probabilities and value functions.
@@ 6 Integration with Reinforcement Learning
where $\text{Proj}_{(\cdot)}$ is operator of taking projection onto $(\cdot)$, and $\mathcal{S}:=\{U:D_{\mathbf{F}}(Z_{t},U)\geq-\alpha\left(\mathbf{F}(Z_{t})-(1-\epsilon)\right)\}$ is the set of safe control actions. We represent this process by
@@
6.1 Integration with Policy Gradient
@@
which is the expected cumulative rewards under policy $\pi_{\theta}$. In the following lemma, we derive the policy gradient calculation for RL with the proposed safety certificate.
@@
$\nabla V^{\theta}=\nabla\int_{\mathcal{I}}R(\tau)\mathbb{P}^{\theta}(\tau)d\tau=\nabla\int_{\bar{\mathcal{I}}}R(\bar{\tau})\mathbb{P}^{\theta}(\bar{\tau})d\bar{\tau}$ (77)
@@
$\nabla V^{\theta}$ $=\nabla\int_{\bar{\mathcal{I}}}R(\bar{\tau})\mathbb{P}^{\theta}(\bar{\tau})d\bar{\tau}=\int_{\bar{\mathcal{I}}}R(\bar{\tau})\nabla\mathbb{P}^{\theta}(\bar{\tau})d\bar{\tau}$ (78)
$=\int_{\bar{\mathcal{I}}}R(\bar{\tau})\mathbb{P}^{\theta}(\bar{\tau})\nabla\log\mathbb{P}^{\theta}(\bar{\tau})d\bar{\tau},$ (79)
where we know that
$\log\mathbb{P}^{\theta}(\bar{\tau})$ $=\nabla\log\biggl[\mathbb{P}(X_{0})\prod_{t=0}^{H_{r}-1}\pi_{\theta}(\hat{U}_{t}\mid X_{t})G(U_{t}\mid X_{t},\hat{U}_{t})$
$\qquad\qquad\qquad\mathbb{P}(X_{t+1}\mid X_{t},U_{t})\biggr]$ (80)
$=\nabla\biggl[\log(\mathbb{P}\left(X_{0}\right))+\sum_{t=0}^{H_{r}-1}\biggl(\log(\pi_{\theta}(\hat{U}_{t}\mid X_{t}))$
$+\log(G(U_{t}\mid X_{t},\hat{U}_{t}))+\log(\mathbb{P}\left(X_{t+1}\mid X_{t},U_{t}\right))\biggr)\biggr]$ (81)
$=\nabla\left[\sum_{t=0}^{H_{r}-1}\log(\pi_{\theta}(\hat{U}_{t}\mid X_{t}))\right]$ (82)
$=\sum_{t=0}^{H_{r}-1}\nabla\log(\pi_{\theta}(\hat{U}_{t}\mid X_{t})).$ (83)
Here, (82) holds because $G$ is independent from the policy $\pi_{\theta}$. Combining (79) and (83), we get (75). ∎
Lemma 2 states that the policy gradients with long-term safety constraints can be computed similarly to standard policy gradients, but with modified control actions and policy distributions. With (75), one can perform policy optimization $\theta\leftarrow\theta+\eta\nabla V^{\theta}$ to train the reinforcement learning agent with $\eta$ being the learning rate [75]. We conclude the procedures to apply the proposed safety filter to policy gradient in Algorithm 4. For each policy gradient iteration in the total $N_{\text{iter}}$ iterations, we collect sample trajectories for $N_{\text{eps}}$ episodes, and calculate the modified policy gradient using (75). Then we can perform policy optimization with the modified policy gradient.
@@
for $i\in\{1,2,3,\cdots,N_{\text{iter}}\}$ do
  for $k\in\{1,2,3,\cdots,N_{\text{eps}}\}$ do
@@
  Estimate $\nabla V^{\theta_{i}}$ in (75) with $\{R(\tau_{k}):k\in\{1,\cdots,N_{\text{eps}}\}\}$
@@
6.2 Integration with Q-learning
@@
= $r(x,u)+\gamma\;{\mathbb{E}}_{{x^{\prime}\sim\mathbb{P}(X_{t+1}\mid X_{t}=x,U_{t}=u)}}\left[\max_{u^{\prime}\in\mathcal{U}}Q\left(x^{\prime},u^{\prime}\right)\right].$
Then, the Q-learning update rule for each iteration $i$ is given by
@@
where $\eta_{i}$ is the learning rate. From [74], when the reward function $r$ is bounded, the learning rate satisfies $0\leq\eta_{i}<1$, and
@@
the Q-learning converges to the optimal Q function asymptotically with probability 1.
The procedures for the proposed safe Q-learning method are summarized in Algorithm 5. Note that the algorithm only learns Q function. There are many choices for how to obtain the control policy using Q function (e.g., $\epsilon$-greedy).